OpenStudy (anonymous):

A kilobyte is 210. A megabyte is 220 bytes. How many kilobytes are in a megabyte? Use the exponent rule for division to find your answer and give your answer in exponential form. Need super help

OpenStudy (anonymous):

Hello Hello Hello

OpenStudy (anonymous):

1 kb = 210 1 Mb = 220 therefpore , no. of Kb in ! Mb = 210 *220 =46200 \[= 4.6200 \times 10^4 =4.62 \times 10^4\]

OpenStudy (anonymous):

Thank you very much and this is the car I like

OpenStudy (anonymous):

Where are you from?

OpenStudy (anonymous):

could you help me?

OpenStudy (anonymous):

Yes I try

OpenStudy (anonymous):

okay What is the solution to g over five equals three over two?